OpenStudy (anonymous):

The sum of three consecutive even numbers is 44 more than the average of these numbers. The product of the smallest and the largest numbers is equal to

OpenStudy (raden):

let that numbers are x, x+2, x+4 according information above, we have an equation : x+x+2+x+4 = (x+x+2+x+4)/3 + 44 3x + 6 = (3x+6)/3 + 44 3x + 6 = x + 2 + 44 3x + 6 = x + 46 solve for x, first